OpenStudy (anonymous):

The table below shows a relationship between x and y. x y -1 6 2 0 5 - 6 8 -12 Write an equation to represent the function relationship presented in the table.

OpenStudy (anonymous):

find the slope between any two points

OpenStudy (anonymous):

them use point slope form y-y1=m(x-x1) where (x1,y1) is any point given to you

OpenStudy (ganpat):

slope = (y2-y1) / (x2-x1) so from given table consider any point as x1 & y1 and next point is x2 & y2.. substitute the value and get m...

OpenStudy (ganpat):

x y -1 6 2 0 5 - 6 8 -12 m = (0-6)/ (2-(-1)) = -6/3 = -2...

OpenStudy (ghazi):

just write the equation of the line using two pairs of point you'll get the relation ... (y-y1)=(y2-y1)/(x2-x1)* (x-x1)

OpenStudy (anonymous):

i got the same answer as Ganpat

OpenStudy (anonymous):

It's a good thing this is a linear relationship :)